OpenStudy (kikuo):

Which is faster? 213 miles/second or 4283 millimeters/hour This is what I have so far. Is it correct? 213 mil/ 1 sec x 60 sec/1 min=12,780 mil/ 1 min x 60 min / 1 hr=766,800 mil/hr

OpenStudy (3mar):

Could I help you?

OpenStudy (kikuo):

Yes, I know this isn't the right answer now, but I'm not sure what I'm doing wrong.

OpenStudy (kikuo):

I first tried to convert the seconds to hours. Then I tried to convert the miles to cm, but I got a huge wrong number.

OpenStudy (kikuo):

This was my final answer 213 ml/ sec x 60 sec/ 1 min = 213x60 ml / 1 min x 60 min / 1 hr = 213 x 60 x 60 ml / hr = 766,800 mi / hr 766,800 ml / hr x 160934 cen / 1 ml = 766,800 x 160934 / hr = 123404191200 / 1 hr

OpenStudy (3mar):

1. Pleasure is mine to help you! 2. I did not judge your answer correct or not, what I am going to do is to show and guide you how to do so... alright?

OpenStudy (kikuo):

Alright

OpenStudy (3mar):

Great! The idea behind such problems is that you convert the unit of one quantity \ amount to the other unit of the other quantity(s)

OpenStudy (kikuo):

I realized one mistake. I was using cm instead of mm.

OpenStudy (kikuo):

213 ml/ sec x 60 sec/ 1 min = 213x60 ml / 1 min x 60 min / 1 hr = 213 x 60 x 60 ml / hr = 766,800 mi / hr 766,800 ml / hr x 1 609 344 mm / 1 ml = 766,800 x1 609 344 / hr = ??? I have this now

OpenStudy (3mar):

As an example of this: